A .360 M solution of KNO3 which contains 125 grams of solute would have a volume of ____

A) 3.60 x103 ML

B) 7.5 X10^3mL

C) 3.43 X 10^3 mL

D) 1.25 x 10^3 mL
Answer #1

Option C

Molarity o) - w to mut 1000 W MW 1600 M V = 125 + 1000 lot I 0-36 v = 3.43 + 10 m2